OpenStudy (anonymous):

S=P/(q+p(1-q)) solve for p

OpenStudy (anonymous):

big p or small p

OpenStudy (anonymous):

\[S=\frac{P}{q+p(1-q)}\]

OpenStudy (abb0t):

\[S = \frac{ P }{ q+p(1-q) }=S(q+p-pq)=P \] distribute the S across the parenthesis, and use algebra

OpenStudy (anonymous):

for small p multiply both sides by q+p(1-q)